# Envy-Free Allocation of Indivisible Goods via Noisy Queries

### v1

@@ 1 Introduction
The fair allocation of indivisible goods (henceforth referred to as “items”) is a fundamental problem at the intersection of computer science and economics with diverse applications, and has been studied extensively in recent years. This topic has been investigated under a wide variety of fairness criteria, each having its own benefits and implications; for some recent surveys, see Moulin, (2019); Aziz, (2020); Walsh, (2020); Suksompong, (2021); Amanatidis et al., (2023). Among these fairness notions, envy-freeness (and its variations) is among the most fundamental and widely-studied. In an envy-free allocation, each agent values the set of items they receive at least as much as they value the items given to any other single agent. Since envy-freeness can be impossible to satisfy (e.g., when there is only one item) or computationally hard to achieve, it is common to study relaxations such as envy-freeness up to one item (EF1).
A prevailing assumption in the literature is that the valuations of the agents are known exactly, or can be queried to learn the exact value. Motivated by real-world settings in which the information gathered is imperfect (e.g., agents may make mistakes in their evaluations, or the evaluations may fluctuate over time), we are interested in scenarios where such queries may be noisy. We focus on the fundamental model of additive Gaussian noise; see Remark 1 in Section 2 for a discussion on our noise model’s justification and limitations. This model of noisy querying matches that of the extensively studied multi-armed bandit problem Lattimore and Szepesvári, (2020), and also falls under the general topic of algorithms with noisy queries (e.g., see Feige et al., (1994); Addanki et al., (2021); Zhu et al., (2023)), though the latter typically involves discrete noise models (e.g., binary-output queries that get flipped with some probability).
@@ 1.1 Related Work
Surveys of these notions, as well as other fairness notions for allocating indivisible goods, can be found, e.g., in Amanatidis et al., (2023).
Envy-free allocation via queries. In most of the existing literature on envy-freeness (and more generally, fair division), it is assumed that the entire set of valuations is known in advance. However, a recent line of works has sought to understand various notions of query complexity. In particular, under additive valuations, noise-free queries on bundles of items were studied in Oh et al., (2021); Bu et al., (2024), with Oh et al., (2021) considering value-based queries and Bu et al., (2024) considering comparison-based queries. In both cases, EF1 was shown to be achievable using a logarithmic number of queries. For the stronger notion of EFX, the query complexity increases to linear under additive valuations Oh et al., (2021) and exponential under more general (non-additive) valuations Plaut and Roughgarden, 2020a .
The preceding works are all fundamentally different from ours due to the queries being noiseless and being applied to bundles of items rather than individual items. Somewhat closer to our work is a recent study of the round-robin algorithm under additive valuations with potentially noisy queries Li et al., (2025). These authors consider comparison-based queries to pairs of items, as well as value-based queries to individual items, giving various upper and lower bounds with at least an $nm$ dependence on the number of agents $n$ and items $m$. The major differences compared to our work are outlined as follows:
@@
  • To make our (more challenging) problem feasible, we introduce a “gap” parameter indicating how negative the (unknown) optimal envy is, i.e., how far the corresponding allocation is from violating EF. Under the assumption of such a gap, we construct algorithms that do not need to query every item. In contrast, the lower bounds in Li et al., (2025) reveal that, as one would expect, implementing the round-robin algorithm exactly requires querying every item at least once.
@@
We note that the above-mentioned comparison-based query model is an instance of dueling bandit feedback Sui et al., (2018), and the additive Gaussian noise model is an instance of regular multi-armed bandit feedback (Lattimore and Szepesvári,, 2020, Ch. 4). We proceed to survey some other (less closely related) works that adopt a multi-armed bandit viewpoint.
@@
Other bandit-based settings involving envy-freeness. In Procaccia et al., (2024), a setting was considered in which items arrive sequentially and are allocated to agents in an online manner. Their goal involves achieving envy-freeness (or proportionality) in expectation, but the main objective itself is a cumulative social welfare measure. Note that in our setup, envy-freeness in expectation could be obtained trivially (with no queries) by simply assigning items uniformly at random. Other differences in their work include only having finitely many “types” of item, and having to allocate items immediately as they arrive (i.e., the online setting). Other related works in the online setting, typically seeking objectives based on Nash welfare, include Sinha et al., (2023); Bhattacharya et al., (2024); Verma et al., (2024); Yamada et al., (2024); Schiffer and Zhang, (2025).
@@
In Peters et al., (2022), a robust rent division problem is studied, with the goal of obtaining envy-free allocations robust to misspecified values. A major difference from our work is that they specifically focus on a “room allocation” problem in which every agent gets one room (i.e., a “matching” is formed). In contrast, we focus on a two-agent problem with arbitrarily many items. At a technical level, their query complexity indicates taking $O\big(\frac{1}{\varepsilon^{2}}\big)$ samples of each (agent, room) pair to attain a probability of envy-freeness within $\varepsilon$ of optimal, whereas in our problem we can often have fewer queries than items.
@@
Repeated allocation problems with bandit feedback. Another line of work on item allocation with bandit feedback has considered scenarios where every round consists of proposing an entire allocation, rather than querying just one item and/or agent. The goal is typically to optimize some long-term measure of fairness. For instance, see Talebi and Proutiere, (2018) for a study of proportional fairness in allocating tasks to servers, Lim et al., (2024) for an egalitarian matching-based assignment problem, and Harada et al., (2025) for a related problem of maximizing the minimum utility aggregated across a long sequence of allocations.
@@
Other fairness notions in bandit algorithms. For other issues of fairness in bandit algorithms (not involving item allocation), we refer the interested reader to Li et al., (2019); Hossain et al., (2021); Patil et al., (2021); Barman et al., (2023); Banihashem et al., (2023); Sawarni et al., (2023); Russo and Vannella, (2024) and the references therein. To name just one example, in Hossain et al., (2021) each arm is valued differently by different agents, and the goal is to identify a distribution over the arms that maximizes the Nash welfare. Since item allocation is central to our work but is not considered in these works, we do not delve into the details.
@@ 1.2 Our Contributions
at least when $\Delta\gg m^{1/4}\log^{2}m$. Moreover, the upper bound is based on a non-adaptive querying strategy and a polynomial-time allocation strategy, whereas the lower bound holds even under adaptive queries and arbitrary computational complexity. We note that the seemingly unconventional $\frac{m^{2.5}}{\Delta^{2}}$ scaling can be more naturally viewed as $\frac{\sqrt{m}}{(\Delta/m)^{2}}$ with $\Delta/m$ representing a “normalized gap”; indeed, quadratic dependencies on gaps are ubiquitous in pure exploration problems for multi-armed bandits Lattimore and Szepesvári, (2020). Some intuition on the $\sqrt{m}$ numerator will be given in Section 4.
@@ 2 Problem Setup
We consider fairly allocating $m$ indivisible items between two agents $a$ and $b$, where the agents have their respective utility $u^{a}_{i}$ and $u^{b}_{i}$ for each item $i$, and $u^{a}_{i},u^{b}_{i}\in[0,1]$. Here the restriction to $[0,1]$ is for convenience, and could be obtained from any fixed interval $u_{\min},u_{\max}]$ by shifting and rescaling. Our problem setup could naturally be extended to more than two agents, but the two-agent setting is a fundamental starting point that already comes with considerable technical challenges. See Section 7 for some further discussion on the $n$-agent scenario.
@@ 3 An Initial Suboptimal Upper Bound
Regarding the first dot point, a key weakness in the naive approach is the reliance on each individual item’s confidence interval when estimating the value of an entire bundle, and implicitly assuming that the errors always accumulate in the worst manner possible (thus multiplying the amount of error by the bundle size).
@@ 4 An Improved Upper Bound
(Assumption on $\Delta$) In general, $\Delta$ lies in the range $[0,m]$, meaning that the restriction to $\Delta\geq m^{1/4}\log^{2}m$ captures “most” of the possible scaling regimes. Nevertheless, it would be of interest to handle $\Delta\leq O(m^{1/4})$ as well. The lower bound on $\Delta$ arises in our analysis for somewhat technical reasons: we want to control a quantity (related to “smoothness” in the second dot point above) to be at most ${\widetilde{O}}\big(\frac{m}{\sqrt{q}}\big)$, which we are only able to achieve if $q\ll m^{2}$. In contrast, attaining the desired bound $q={\widetilde{O}}\big(\frac{m^{2.5}}{\Delta^{2}}\big)$ when $\Delta\ll m^{1/4}$ requires us to have $q\gg m^{2}$. Overall, while we do not wish to confidently claim anything around this discussion, we expect that precluding the regime $\Delta\leq O(m^{1/4})$ is not merely an artifact/weakness in our analysis, but rather, that this regime requires fundamentally different algorithms (e.g., based on actually forming bundles rather than using the item-by-item approach of (4)).
@@ 6 General Noise Levels
  • The case $q<m$ is handled in Theorem 102 in Appendix C.4, which states that if (i) $\Delta^{2}>160\sigma m^{3/2}\log^{2}m$, (ii) $\Delta^{4}>160^{2}m^{3}\sigma^{4}\log^{2}m$, and (iii) $\Delta<2\sigma^{2}m$, then a sufficient number of queries is
@@ 7 Conclusion
Finally, regarding other fairness notions, we comment on another well-known fairness notion, proportionality (e.g., see Amanatidis et al., (2023)). An allocation is said to be proportional if it gives each agent a utility least as high as the agent’s proportional share, defined as $1/n$ times their utility for the set of all items. In the case of $n=2$ agents and additive utilities, envy-freeness and proportionality are equivalent: A direct comparison of the fairness definitions yields that an agent has an envy of $r>0$ (resp., $r<0$) if and only if the agent’s utility is below (resp., above) their proportional share by $r/2$. Therefore, all of our results can be directly transferred to proportionality. As with envy-freeness, extending these results to $n$ agents for proportionality is an interesting direction for future work.
@@ Acknowledgment
This work is supported by the Singapore National Research Foundation (NRF) under its AI Visiting Professorship programme, by the Singapore Ministry of Education under grant number MOE-T2EP20221-0001, and by an NUS Start-up Grant.
@@ Appendix A Probabilistic Tools
(Chernoff bound for Gaussian RVs (Vershynin,, 2018, Ch. 2)) Let $y_{1},\dots,y_{T}$ be independent random variables such that $y_{t}\sim N(\mu,\sigma^{2})$. Then, it holds for any $\varepsilon>0$ that
@@
(Bernstein’s inequality for bounded RVs (Vershynin,, 2018, Ch. 2)) Let $y_{1},\dots,y_{T}$ be independent and identically distributed zero-mean random variables satisfying $|y_{i}|\leq b$ almost surely, and $\mathrm{Var}[y_{i}]=\sigma_{y}^{2}$. Then, we have
@@
(Berry–Esseen theorem for non-identical RVs (Petrov,, 1995, Thm. 5.6)) Let $y_{1},\dots,y_{T}$ be independent random variables with each $y_{t}$ having mean $\mu_{t}$, variance $\sigma_{t}^{2}$, and finite third absolute moment. Define
@@
$\sup_{\gamma\in{\mathbb{R}}}\big|{\mathbb{P}}[Z_{T}\leq\gamma]-{\mathbb{P}}[N(0,1)\leq\gamma]\big|\leq\frac{C\Psi_{T}}{V_{T}^{3/2}}.$ (11)
@@
(Assouad’s lemma (Duchi,, 2023, Sec. 9.5)) Let $\Theta=\{0,1\}^{d}$ for some integer $d\geq 1$, let $\{P_{\theta}:\theta\in\Theta\}$ be a family of probability measures taking values on some space $\mathcal{Y}$, and let $\hat{\theta}:\mathcal{Y}\to\Theta$ be any estimator of $\Theta$ based on an outcome in $\mathcal{Y}$. Define the Hamming distance $d_{H}(\theta,\theta^{\prime})=\sum_{i=1}^{d}\mathbf{1}\{\theta_{i}\neq\theta^{\prime}_{i}\}$, and the optimal expected risk under a uniform prior as [^5] In (Duchi,, 2023, Thm. 9.5.2) the lemma is stated in terms of the minimax risk instead (i.e., with $\max_{\theta}$ in place of $\frac{1}{|\Theta|}\sum_{\theta\in\Theta}$), but the proof in (Duchi,, 2023, Sec. 9.6.3) is based on lower bounding the minimax risk by the average risk.
@@
(Convexity of KL divergence (Polyanskiy and Wu,, 2025, Thm. 5.1)) The KL divergence $D(P\|Q)$ is jointly convex in its arguments, thus implying the following via Jensen’s inequality: Let $\{P_{j}\}$ and $\{Q_{j}\}$ be two countable [^6] The counterpart with continuous indices also holds, but we will only need the discrete version, and it is more convenient to state in this form. collections of distributions on a common alphabet, let $\lambda$ be an arbitrary distribution over the indices $\{j\}$ (i.e., the values $\lambda_{j}$ are non-negative and sum to $1$), and define the mixture distributions $P(x)=\sum_{j}\lambda_{j}P_{j}(x)$ and $Q(x)=\sum_{j}\lambda_{j}Q_{j}(x)$. Then, we have
@@
(Consequence of chain rule for KL divergence, e.g., Ex. 15.8 and Eq. (38.22) in Lattimore and Szepesvári, (2020)) Fix positive integers $k$ and $n$. Consider any distributions $P_{1},P^{\prime}_{1},\dotsc,P_{k},P^{\prime}_{k}$ on a common alphabet, and write $P=(P_{1},\dotsc,P_{k})$ and $P^{\prime}=(P^{\prime}_{1},\dotsc,P^{\prime}_{k})$. Then, consider an arbitrary (possibly adaptive) algorithm that, at each time indexed by $t=1,\dotsc,n$, queries an index $i_{t}\in\{1,\dotsc,k\}$ and observes a random sample $y_{t}$ from the $i_{t}$-th distribution (i.e., from $P_{i_{t}}$ under $P$, or from $P^{\prime}_{i_{t}}$ under $P^{\prime}$). Let $Y=(y_{1},\dotsc,y_{n})$ be the resulting sequence of outcomes, and let $P_{Y}$ and $P^{\prime}_{Y}$ be the corresponding joint distributions on these outcomes under $P$ and $P^{\prime}$. Then, it holds that

### v2

@@ 1 Introduction
The fair allocation of indivisible goods (henceforth referred to as “items”) is a fundamental problem at the intersection of computer science and economics with diverse applications, and has been studied extensively in recent years. This topic has been investigated under a wide variety of fairness criteria, each having its own benefits and implications; for some recent surveys, see (Moulin, 2019; Aziz, 2020; Walsh, 2020; Suksompong, 2021; Amanatidis et al., 2023). Among these fairness notions, envy-freeness (and its variations) is among the most fundamental and widely-studied. In an envy-free allocation, each agent values the set of items they receive at least as much as they value the items given to any other single agent. Since envy-freeness can be impossible to satisfy (e.g., when there is only one item) or computationally hard to achieve, it is common to study relaxations such as envy-freeness up to one item (EF1).
A prevailing assumption in the literature is that the valuations of the agents are known exactly, or can be queried to learn the exact value. In this paper, we are interested in scenarios where such queries may be noisy, motivated by scenarios such as the following:
  • (i)
    The “utilities” are not measurable or human-chosen, but instead need to be estimated via simulations that are stochastic in nature (e.g., estimating how much a particular resource will help each team in a company’s internal division problem).
  • (ii)
    Each so-called “agent” is actually a group of people, and the “utility” is considered to be an average over the group members. To query an item, we may take a random group member and ask for their valuation. [^1] This formulation assumes that fairness is only required on an averaged group level, not an individual level.
  • (iii)
    More generally, noise may arise in settings where agents make mistakes in their evaluations, the evaluations fluctuate over time, and so on.
It is likely highly challenging to devise a general noise model capturing all such scenarios, so as a natural starting point, we focus on the fundamental model of additive Gaussian noise; see Remark 1 in Section 2 for a discussion on our noise model’s justification and limitations. This model of noisy querying matches that of the extensively studied multi-armed bandit problem (Lattimore and Szepesvári, 2020), and also falls under the general topic of algorithms with noisy queries (e.g., see (Feige et al., 1994; Addanki et al., 2021; Zhu et al., 2023)), though the latter typically involves discrete noise models (e.g., binary-output queries that get flipped with some probability).
@@ 1.1 Related Work
Surveys of these notions, as well as other fairness notions for allocating indivisible goods, can be found, e.g., in (Amanatidis et al., 2023).
Envy-free allocation via queries. In most of the existing literature on envy-freeness (and more generally, fair division), it is assumed that the entire set of valuations is known in advance. However, a recent line of works has sought to understand various notions of query complexity. In particular, under additive valuations, noise-free queries on bundles of items were studied in (Oh et al., 2021; Bu et al., 2024), with (Oh et al., 2021) considering value-based queries and (Bu et al., 2024) considering comparison-based queries. In both cases, EF1 was shown to be achievable using a logarithmic number of queries. For the stronger notion of EFX, the query complexity increases to linear under additive valuations (Oh et al., 2021) and exponential under more general (non-additive) valuations (Plaut and Roughgarden, 2020).
The preceding works are all fundamentally different from ours due to the queries being noiseless and being applied to bundles of items rather than individual items. Somewhat closer to our work is a recent study of the round-robin algorithm under additive valuations with potentially noisy queries (Li et al., 2025). These authors consider comparison-based queries to pairs of items, as well as value-based queries to individual items, giving various upper and lower bounds with at least an $nm$ dependence on the number of agents $n$ and items $m$. The major differences compared to our work are outlined as follows:
@@
  • To make our (more challenging) problem feasible, we introduce a “gap” parameter indicating how negative the (unknown) optimal envy is, i.e., how far the corresponding allocation is from violating EF. Under the assumption of such a gap, we construct algorithms that do not need to query every item. In contrast, the lower bounds in (Li et al., 2025) reveal that, as one would expect, implementing the round-robin algorithm exactly requires querying every item at least once.
@@
We note that the above-mentioned comparison-based query model is an instance of dueling bandit feedback (Sui et al., 2018), and the additive Gaussian noise model is an instance of regular multi-armed bandit feedback ([)Ch. 4]lattimore2020bandit. We proceed to survey some other (less closely related) works that adopt a multi-armed bandit viewpoint.
@@
Other bandit-based settings involving envy-freeness. In (Procaccia et al., 2024), a setting was considered in which items arrive sequentially and are allocated to agents in an online manner. Their goal involves achieving envy-freeness (or proportionality) in expectation, but the main objective itself is a cumulative social welfare measure. Note that in our setup, envy-freeness in expectation could be obtained trivially (with no queries) by simply assigning items uniformly at random. Other differences in their work include only having finitely many “types” of item, and having to allocate items immediately as they arrive (i.e., the online setting). Other related works in the online setting, typically seeking objectives based on Nash welfare, include (Sinha et al., 2023; Bhattacharya et al., 2024; Yamada et al., 2024; Schiffer and Zhang, 2025; Verma et al., 2026).
@@
In (Peters et al., 2022), a robust rent division problem is studied, with the goal of obtaining envy-free allocations robust to misspecified values. A major difference from our work is that they specifically focus on a “room allocation” problem in which every agent gets one room (i.e., a “matching” is formed). In contrast, we focus on a two-agent problem with arbitrarily many items. At a technical level, their query complexity indicates taking $O\big(\frac{1}{\varepsilon^{2}}\big)$ samples of each (agent, room) pair to attain a probability of envy-freeness within $\varepsilon$ of optimal, whereas in our problem we can often have fewer queries than items.
@@
Repeated allocation problems with bandit feedback. Another line of work on item allocation with bandit feedback has considered scenarios where every round consists of proposing an entire allocation, rather than querying just one item and/or agent. The goal is typically to optimize some long-term measure of fairness. For instance, see (Talebi and Proutiere, 2018) for a study of proportional fairness in allocating tasks to servers, (Lim et al., 2024) for an egalitarian matching-based assignment problem, and (Harada et al., 2025) for a related problem of maximizing the minimum utility aggregated across a long sequence of allocations.
@@
Other fairness notions in bandit algorithms. For other issues of fairness in bandit algorithms (not involving item allocation), we refer the interested reader to (Li et al., 2019; Hossain et al., 2021; Patil et al., 2021; Banihashem et al., 2023; Barman et al., 2023; Sawarni et al., 2023; Russo and Vannella, 2024) and the references therein. To name just one example, in (Hossain et al., 2021) each arm is valued differently by different agents, and the goal is to identify a distribution over the arms that maximizes the Nash welfare. Since item allocation is central to our work but is not considered in these works, we do not delve into the details.
@@ 1.2 Our Contributions
at least when $\Delta\gg m^{1/4}\log^{2}m$. Moreover, the upper bound is based on a non-adaptive querying strategy and a polynomial-time allocation strategy, whereas the lower bound holds even under adaptive queries and arbitrary computational complexity. We note that the seemingly unconventional $\frac{m^{2.5}}{\Delta^{2}}$ scaling can be more naturally viewed as $\frac{\sqrt{m}}{(\Delta/m)^{2}}$ with $\Delta/m$ representing a “normalized gap”; indeed, quadratic dependencies on gaps are ubiquitous in pure exploration problems for multi-armed bandits (Lattimore and Szepesvári, 2020). Some intuition on the $\sqrt{m}$ numerator will be given in Section 4.
@@ 2 Problem Setup
We consider fairly allocating $m$ indivisible items between two agents $a$ and $b$, where the agents have their respective utility $u^{a}_{i}$ and $u^{b}_{i}$ for each item $i$, and $u^{a}_{i},u^{b}_{i}\in[0,1]$. Here the restriction to $[0,1]$ is for convenience, and could be obtained from any interval of the form $0,u_{\max}]$ by rescaling. Our problem setup could naturally be extended to more than two agents, but the two-agent setting is a fundamental starting point that already comes with considerable technical challenges. See Section 7 for some further discussion on the $n$-agent scenario.
@@ 3 An Initial Suboptimal Upper Bound
Regarding the first dot point, a key weakness in the naive approach is the reliance on each individual item’s confidence interval when estimating the value of an entire bundle, and implicitly assuming that the errors always accumulate in the worst manner possible (thus multiplying the amount of error by the bundle size). Note that this weakness concerns the analysis rather than the algorithm itself, so it is conceivable that Algorithm˜1 could satisfy stronger guarantees if analyzed more carefully.
@@ 4 An Improved Upper Bound
(Assumption on $\Delta$) In general, $\Delta$ lies in the range $[0,m]$, meaning that the restriction to $\Delta\geq m^{1/4}\log^{2}m$ captures “most” of the possible scaling regimes. Nevertheless, it would be of interest to handle $\Delta\leq O(m^{1/4})$ as well. The lower bound on $\Delta$ arises in our analysis for somewhat technical reasons: we want to control a quantity (related to “smoothness” in the second dot point above) to be at most ${\widetilde{O}}\big(\frac{m}{\sqrt{q}}\big)$, which we are only able to achieve if $q\ll m^{2}$. In contrast, attaining the desired bound $q={\widetilde{O}}\big(\frac{m^{2.5}}{\Delta^{2}}\big)$ when $\Delta\ll m^{1/4}$ requires us to have $q\gg m^{2}$. We also note that certain regimes of sufficiently small $\Delta$ may have computational barriers in view of the fact finding an envy-free allocation (i.e., $\Delta=0$) is NP-hard due to a reduction from PARTITION (Lipton et al., 2004).
Overall, while we do not wish to confidently claim anything around this discussion, we expect that precluding the regime $\Delta\leq O(m^{1/4})$ is not merely an artifact/weakness in our analysis, but rather, that this regime requires fundamentally different algorithms (e.g., based on actually forming bundles rather than using the item-by-item approach of (4)).
@@ 6 General Noise Levels
  • The case $q<m$ is handled in Theorem 102 in Appendix C.4, which states that if (i) $\Delta^{2}>160\sigma m^{3/2}\log^{2}m$, (ii) $\Delta^{4}>160^{2}m^{3}\sigma^{4}\log^{2}m$, and (iii) $\Delta<\min\{2\sigma^{2}m,\sqrt{2}\sigma m\}$, then a sufficient number of queries is
@@ 7 Conclusion
Finally, regarding other fairness notions, we comment on another well-known fairness notion, proportionality (e.g., see (Amanatidis et al., 2023)). An allocation is said to be proportional if it gives each agent a utility least as high as the agent’s proportional share, defined as $1/n$ times their utility for the set of all items. In the case of $n=2$ agents and additive utilities, envy-freeness and proportionality are equivalent: A direct comparison of the fairness definitions yields that an agent has an envy of $r>0$ (resp., $r<0$) if and only if the agent’s utility is below (resp., above) their proportional share by $r/2$. Therefore, all of our results can be directly transferred to proportionality. As with envy-freeness, extending these results to $n$ agents for proportionality is an interesting direction for future work.
@@ Acknowledgment
Y.L. is supported by NTU Start-up Grant #025311-00001. J.S. is supported by the Singapore National Research Foundation (NRF) under its AI Visiting Professorship programme. W.S. is supported by the Singapore Ministry of Education under grant number MOE-T2EP20221-0001 and by an NUS Start-up Grant.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Probabilistic Tools
(Chernoff bound for Gaussian RVs ([)Ch. 2]Ver18) Let $y_{1},\dots,y_{T}$ be independent random variables such that $y_{t}\sim N(\mu,\sigma^{2})$. Then, it holds for any $\varepsilon>0$ that
@@
(Bernstein’s inequality for bounded RVs ([)Ch. 2]Ver18) Let $y_{1},\dots,y_{T}$ be independent and identically distributed zero-mean random variables satisfying $|y_{i}|\leq b$ almost surely, and $\mathrm{Var}[y_{i}]=\sigma_{y}^{2}$. Then, we have
@@
(Berry–Esseen theorem for non-identical RVs ([)Thm. 5.6]Petrov1995) Let $y_{1},\dots,y_{T}$ be independent random variables with each $y_{t}$ having mean $\mu_{t}$, variance $\sigma_{t}^{2}$, and finite third absolute moment. Define
@@
$\sup_{\gamma\in{\mathbb{R}}}\big|{\mathbb{P}}[Z_{T}\leq\gamma]-{\mathbb{P}}[N(0,1)\leq\gamma]\big|\leq\frac{C\Psi_{T}}{V_{T}^{3/2}}$ (11)
@@
(Assouad’s lemma ([)Sec. 9.5]duchi2016lecture) Let $\Theta=\{0,1\}^{d}$ for some integer $d\geq 1$, let $\{P_{\theta}:\theta\in\Theta\}$ be a family of probability measures taking values on some space $\mathcal{Y}$, and let $\hat{\theta}:\mathcal{Y}\to\Theta$ be any estimator of $\Theta$ based on an outcome in $\mathcal{Y}$. Define the Hamming distance $d_{H}(\theta,\theta^{\prime})=\sum_{i=1}^{d}\mathbf{1}\{\theta_{i}\neq\theta^{\prime}_{i}\}$, and the optimal expected risk under a uniform prior as [^6] In ([)Thm. 9.5.2]duchi2016lecture the lemma is stated in terms of the minimax risk instead (i.e., with $\max_{\theta}$ in place of $\frac{1}{|\Theta|}\sum_{\theta\in\Theta}$), but the proof in ([)Sec. 9.6.3]duchi2016lecture is based on lower bounding the minimax risk by the average risk.
@@
(Convexity of KL divergence ([)Thm. 5.1]polyanskiy2025information) The KL divergence $D(P\|Q)$ is jointly convex in its arguments, thus implying the following via Jensen’s inequality: Let $\{P_{j}\}$ and $\{Q_{j}\}$ be two countable [^7] The counterpart with continuous indices also holds, but we will only need the discrete version, and it is more convenient to state in this form. collections of distributions on a common alphabet, let $\lambda$ be an arbitrary distribution over the indices $\{j\}$ (i.e., the values $\lambda_{j}$ are non-negative and sum to $1$), and define the mixture distributions $P(x)=\sum_{j}\lambda_{j}P_{j}(x)$ and $Q(x)=\sum_{j}\lambda_{j}Q_{j}(x)$. Then, we have
@@
(Consequence of chain rule for KL divergence, e.g., Ex. 15.8 and Eq. (38.22) in (Lattimore and Szepesvári, 2020)) Fix positive integers $k$ and $n$. Consider any distributions $P_{1},P^{\prime}_{1},\dotsc,P_{k},P^{\prime}_{k}$ on a common alphabet, and write $P=(P_{1},\dotsc,P_{k})$ and $P^{\prime}=(P^{\prime}_{1},\dotsc,P^{\prime}_{k})$. Then, consider an arbitrary (possibly adaptive) algorithm that, at each time indexed by $t=1,\dotsc,n$, queries an index $i_{t}\in\{1,\dotsc,k\}$ and observes a random sample $y_{t}$ from the $i_{t}$-th distribution (i.e., from $P_{i_{t}}$ under $P$, or from $P^{\prime}_{i_{t}}$ under $P^{\prime}$). Let $Y=(y_{1},\dotsc,y_{n})$ be the resulting sequence of outcomes, and let $P_{Y}$ and $P^{\prime}_{Y}$ be the corresponding joint distributions on these outcomes under $P$ and $P^{\prime}$. Then, it holds that